OpenStudy (anonymous):

Which is the slope and y-intercept of this graph? A. slope = -1; y-intercept = 4 B. slope = 1; y-intercept = 4 C. slope = -4; y-intercept = 4 D. slope = -1; y-intercept = 0

OpenStudy (anonymous):

Where's the graph?

OpenStudy (anonymous):

o sory one min

OpenStudy (anonymous):

open with paint

OpenStudy (anonymous):

y=mx+b. Pick two coordinate points, (0,4) and (4,0). The y intercept is where the line crosses the y-axis. Here it's 4. Using the slope equation do y2-y1 / (x2-x1). Your answer should be y = -x+4.

OpenStudy (anonymous):

so what would the letter be

OpenStudy (anonymous):

so it would be A

OpenStudy (anonymous):

ok yhx